OpenStudy (anonymous):

A total of 1000 was invested, part of it at 12% and part at 10%. The total yield was $116. Find the amount invested at each rate.

OpenStudy (anonymous):

Solve the following for x and y:\[\left\{x+y=1000, \frac{12}{100}x+\frac{10}{100}y=116\right\}\]\[{x = 800, y = 200}\]

OpenStudy (anonymous):

Thank you for the medal.